OpenStudy (anonymous):

Medal A pharmaceutical company combined 485,100 mg of acetaminophen with 9,900 mg of oxycodone to fill an order from a pharmacy for a powerful prescription pain-killer. The pain-killers come in 550 mg tablets. How many tablets were created by the company for this shipment? ? tablets

OpenStudy (anonymous):

i think the answer 900

OpenStudy (anonymous):

485100 + 9900 = 495090 495090/552

OpenStudy (anonymous):

550*

OpenStudy (yolo_king):

= 485100 + 9900 = 495,000 495000/550 = 900 tablets So they produced 900 tablets You are right!

OpenStudy (anonymous):

:)